Circular No. 6373

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) C. Lisse, M. Mumma, and R. Petre, NASA Goddard Space Flight

Center; and K. Dennerl, J. Schmitt, J. Englhauser, and J. Truemper, Max-Planck-Institut für Extraterrestrische Physik, Garching, report the discovery of x-rays from comet C/1996 B2 with ROSAT — the first such detections from any comet: “The observations took place on nine different occasions during Mar. 26–28 (see IAUC 6350). The brightest parts of the comet’s x-ray image are diffuse but crescent shaped, offset sunward by ∼ 6′ from the nucleus (projected distance ∼ 30 000 km), and extend to about ± 8′ (± 40 000 km) in the direction perpendicular to the sun-comet direction. The observed radiation shows a clear x-ray (rather than ultraviolet) signature and is strongly time variable on the order of hours. An image can be

found under http://www.rosat.mpe-garching.mpg.de/~jer/rda/comet

/hyakutake.html. Preliminary estimates yield peak surface brightness countrates of ∼ 0.01 count s^{−1} arcmin^{−2}, with the comet being at Δ = 0.13 AU. A probable mechanism for the observed radiation is scattering of solar x-rays by material in the comet’s coma. Another possibility is that the radiation is derived from energy deposition by the solar wind.”

COMET C/1996 B1 (SZCZEPANSKI)

Ephemeris extension from orbital elements on MPC 26879:

1996 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Mar. 28 9^{h}34.^{m}64 − 1^{o}07.^{′}5 0.733 1.612 136.^{o}6 25.^{o}2  8.9
Apr. 2 9 28.26 − 4 55.3 0.806 1.643 130.7 27.5  9.2
7 9 24.06 − 8 04.5 0.885 1.676 125.4 29.1  9.5
12 9 21.64 −10 42.6 0.968 1.712 120.6 30.3  9.8
17 9 20.70 −12 56.4 1.055 1.749 116.3 31.0 10.0
22 9 21.00 −14 51.0 1.144 1.787 112.4 31.3 10.3
27 9 22.33 −16 31.0 1.235 1.828 108.8 31.4 10.6
May 2 9 24.53 −17 59.5 1.328 1.869 105.5 31.3 10.8
7 9 27.45 −19 19.2 1.421 1.911 102.4 31.0 11.1
12 9 31.00 −20 32.0 1.516 1.955  99.5 30.6 11.3
17 9 35.08 −21 39.5 1.611 1.999  96.7 30.2 11.5
22 9 39.64 −22 43.0 1.706 2.045  94.1 29.6 11.8
27 9 44.60 −23 43.6 1.801 2.091  91.6 29.0 12.0
June 1 9 49.91 −24 41.9 1.897 2.137  89.1 28.3 12.2
6 9 55.53 −25 38.6 1.992 2.185  86.8 27.6 12.4
1996 April 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green